The range of human hearing extends from approximately 20 Hz to 20,000 Hz. Find the wavelengths of these extremes at a temperature of 27°C


1
Expert's answer
2022-02-23T08:30:46-0500

The wavelengths of a sound wave in the air

"\lambda=\frac{v}{f}"

So

"\lambda_{\max}=\rm\frac{340\; m/s}{20\: Hz}=17\: m"

"\lambda_{\min}=\rm\frac{340\; m/s}{20000\: Hz}=17\: mm"